3. Useful formulas: fx = MxN a. Find frictional force and acceleration in the case given below. F = 20 N 4 kg Hs = 0.6,HR = 0 Find frictional force
Answer #1

QUESTION 3 (PART a)

7 Drawing the given Scenario Solving the problem. • given, mass= 4 kg on N F -20N lls = 0.6 lk = 0.4 F-20N E surface Wong 99-

QUESTION 3 (PART b)

→ Again drawing the given scenario & solving the problem. • Resolving the force F = 10V2N in component along % 84 axi. AFY AN(Frnet), = man = ma= Fa- fx = 10- 8.76 a 10-8.76 = 1.24 - 0-31 m/s? m4s So, answers are N= 24-2 N fx= 8.760 first time na maj